OpenStudy (anonymous):

Please help:) I forgot how to solve a part of the equation, that will be provided below. Thanks!!

OpenStudy (anonymous):

 I forgot how to solve that portion of the equation

OpenStudy (eli_moses):

use the distributive property do you know what it is?

OpenStudy (anonymous):

Yes, but the fraction confuses me

OpenStudy (eli_moses):

hmm... ok so lets start with multiplying -1/3 to x remember that the x has an invisible 1 next to it so basically your multiplying -1/3 to 1x

OpenStudy (anonymous):

Okay, so multiplying that would give me -1 ?

OpenStudy (anonymous):

- 1/3 * 1/1 (multiply 1/1 by 3) = -3/3 = -1 would that be correct?

OpenStudy (eli_moses):

no if you multiply a number to 1, it gives you the same number like 1 x 3 is 3 not 1

OpenStudy (anonymous):

Yeah, I know that but wouldn't you have to do this > or am i wrong?